Fermi National Laboratory


Accelerator Update for February 11, 2005 - February 14, 2005

Friday February 11
The day shift began with Linac and Booster running beam.

Cryo system technicians reported at 9:38 AM that they were warming up the Tevatron (TeV) sector D1 to change a Kautzky valve.

Operations received permission at 9:42 AM to rack in the 13.8 kV power for the Main Injector (MI).

Operations established beam to MiniBooNE at 12:41 PM.

At 2:51 AM, an Accumulator Septa power supply (A:ISEP2V) tripped off. It reset okay.

Operators began turning on the Antiproton Source (Pbar) at 2:54 PM.

Meson area MCenter experimenters asked for beam at 3:39 PM.

Operations reported that they had the Collider Detector at Fermilab (CDF) interlocked by 3:54 PM.

An RF expert reported at 3:55 PM that he was turning on the RF for MI and the TeV.

Operations received permission to run beam in all machines at 5:31 PM.

Operators had the D-Zero (D0) collision hall interlocked by 6:57 PM.

Operations reported at 8:18 PM that they had the MI beam permit.

A water tech reported at 9:18 PM that the Meson LCW system was operating. The power supplies could be turned on.

Operators reported at 9:26 PM that they had search and secured the TeV.

Operations reported that they had established beam to the Accumulator at 9:49 PM.

A TeV expert reported at 11:56 PM that he had found a problem with a TeV separator (B17) cable. Another expert came in to help.

Saturday February 12
The midnight shift began with the TeV in access to work on the B17 separator, with Pbar stacking slowly, and with MiniBooNE taking beam.

A Pbar expert came in and at 1:55 AM tuned the machine to an efficient antiproton production.

The TeV experts reported at 4:24 AM that they had completed the B17 separator repairs.

At 5:04 AM, Pbar's A:ISEP2V tripped off again. It then tripped off three more times in the next hour. An expert was called in to work on it.

Operations reported at 8:38 AM that the TeV checkout was completed.

The expert working on A:ISEP2V reported at 8:47 AM that the repair he was beginning would hold off stacking. He completed his work and stacking resumed at 9:16 AM.

At 12:54 PM, a Booster RF station (BRF15) tripped off. Operator investigated and then called in an expert.

At 2:25 PM, the Linac, MI, and TeV machines had a number of devices all trip off at the same time, as if there was a power glitch.

Stacking resumed at 3:12 PM.

Operators had all the Meson enclosures and experimental collision halls interlocked by 4:15 PM. Operations established beam to MCenter and then turned the beam off so the experimenters could access their expermental hall.

Operations established beam to MTest around 6:32 PM.

Operations put the TeV through a wet squeeze at 8:31 PM.

Operations put the TeV into shot setup around 9:20 PM.

Sunday February 13
The midnight shift began with the TeV in shot setup, with Switchyard (SY) beam off due to shot setup, and with MiniBooNE taking beam.

Operations established store 3977 at 12:56 AM with an initial luminosity of 69.4E30. Pbar resumed stacking and MCenter and MTest resumed taking beam.

Operations put the Recycler into shot setup to stash antiprotons 5:53 AM. This shut off SY beam.

Stashing ended by 7:03 AM and stacking resumed.

At 7:26 AM, Operations reported that they were having trouble getting beam to MCenter. MTest was content with the beam they were getting. Operators contacted a SY expert.

At 10:08 AM, a D0 collision hall door opened, which caused the interlocks to drop and store 3977 to abort. The TeV quenched at sector D1. Later, operators determined that the door was not latched properly and a slight air pressure changed had probably caused the door to move enough to trip the interlock switch.

Operators prepared CDF and D0 for accesses at 10:15 AM and the TeV for access at 10:56 AM.

CDF and D0 came out of access and Operations put the TeV through a dry squeeze at 3:26 PM.

Operations put the TeV into shot setup at 7:01 PM.

Operations established store 3980 at 10:38 PM with an initial luminosity of 76.2E30. Pbar resumed stacking.

Operations established beam to MCenter and MTest at 11:01 PM.

Monday February 14
The midnight shift began with Operations monitoring store 3980, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

A little after midnight a Pbar expert came in to work on increasing stacking rates.

At 2:54 AM, the MTest beam intensity dropped. Operators investigated.

Operations reported around 3 AM that the Pbar expert had increased the stacking rate to an acceptable level.

At 3:27 AM, MTest experimenters reported that their beam went away. Operators continued their investigation.

At 3:59 AM, Operations reported that a Booster RF station (BRF4) was repeatedly tripping on a power supply overload. An expert was called in.

Operations reported at 4:52 AM that they hadn't found the cause for MTest's loss of beam.

Operations began stashing antiprotons into the Recycler at 6:21 AM. Recycler shots were completed by 8:12 AM.

Stacking resumed at 8:26 AM.

The Plan for Monday
The plan for today is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/14/2005   email Fermilab